OpenStudy (lizz123):

HELP!!! MEDAL!! CREATE A MODEL (LINEAR EQUATION) AND SOLVE! In a recent basketball game between the Storm and the Chargers the numbers of points scored by the Storm in the first half was equal to the number of points scored by the Chargers in the second half of the game. In the first half the Storm scored one-fourth the total points scored in the game. During the second half the Chargers scored 16 points less than one-third the points scored in the game. How many total points were scored during the game?

OpenStudy (lizz123):

@sammixboo

OpenStudy (lizz123):

@dumbcow

OpenStudy (lizz123):

@Kat_Caca_1996

OpenStudy (dumbcow):

Let s and c represent the points scored for each team so total = s+c further we need to tell difference between 1st half and 2nd half points lets use s1,s2 and c1,c2 s = s1+s2 c = c1 +c2 we are given: s1 = c2 s1 = 1/4(s+c) c2 = 1/3(s+c) - 16 we can set s1 and c2 equal ---> 1/4(s+c) = 1/3(s+c) - 16 solve for "s+c" , treat it like a single variable

OpenStudy (lizz123):

isn't s 1/4

OpenStudy (lizz123):

I am confused

OpenStudy (lizz123):

C=16-1/3

OpenStudy (lizz123):

am i correct

OpenStudy (dumbcow):

no change the letters..... s+c = total = t \[\frac{1}{4}t = \frac{1}{3} t - 16\]

OpenStudy (lizz123):

ohh

OpenStudy (lizz123):

and what do i do next

OpenStudy (lizz123):

wait does t=-192

OpenStudy (dumbcow):

solve for t using algebra get t terms on same side \[\frac{1}{3} t - \frac{1}{4}t = 16\]

OpenStudy (dumbcow):

yep well close , you cant have negative points

OpenStudy (lizz123):

so wht do i do next

OpenStudy (dumbcow):

how did you get 192 , it is correct

OpenStudy (lizz123):

1. I multiplied both sides by 12 I moved all terms to one side I Simplified 3t−4t+192 to −t+192 Subtract 192 from both sides Multiply both sides by −1

OpenStudy (lizz123):

that is what i did

OpenStudy (dumbcow):

very good, that is correct way to solve for t

OpenStudy (lizz123):

thnx